The position of a particle is given by ? = 4.50 ?−0.30?, where x has units of meters. What is (a) the average velocity between 2.00 and 3.00 s; (b) the instantaneous velocity at 2.00 s; (c) the instantaneous acceleration at 2.00 s.
Answer #1

Given that,

x = 4.50 e-0.3t

(a)

average velocity between 2.00 and 3.00 s,

vavg = xf - xi / tf - ti

vavg = 4.50 (e-0.3*3 - e-0.30*2)/ (3 - 2)

vavg = -0.64 m/s

(b)

instantaneous velocity at 2.00 s,,

v = dx / dt = d(4.50 e-0.3t) / dt

v = -0.30*4.50*e-0.3*2

v = -0.7408 m/s

(c)

instantaneous acceleration at 2.00 s,

a = dv / dt = d(-0.30*4.50*e-0.3*t) / dt

a = (-0.30)2 * 4.50*e-0.3*2

a = 0.222 m/s^2